OpenStudy (anonymous):

Simplify: 2/5cd - 3/10cd

OpenStudy (anonymous):

\[2/5cd-3/10cd=4/10cd-3/10cd=1/10cd\]

OpenStudy (anonymous):

Can you explain how you got your answer?

OpenStudy (anonymous):

Multiplied the denominator and numerator of 2/5cd by 2 to get a common denominator.

OpenStudy (anonymous):

if youtake the 2/5...it = 4/10 whe you make the Denominater equal